OpenStudy (anonymous):

what is the average value of y = sin2x over the interval ( pi/4, pi/3)

ganeshie8 (ganeshie8):

average value of \(f(x)\) in interval \((a, b)\) = \(\large \mathbb{\frac{1}{b-a}\int_a^b f(x) dx}\)

ganeshie8 (ganeshie8):

take the definite integral in given interval, divide by the difference b-a

OpenStudy (anonymous):

I got, 1/pi/12 ( sin 2x)

OpenStudy (anonymous):

but then I did the integral for ( sin2x) and got - cos2x

ganeshie8 (ganeshie8):

yes evaluate the limits

OpenStudy (anonymous):

-1/pi/6?

OpenStudy (anonymous):

i got that answer but the correct answer is - 1/6pi...

ganeshie8 (ganeshie8):

average value of \(\sin 2x\) in interval \((\pi/4, \pi/3)\) = \(\large \mathbb{\frac{1}{\pi/3-\pi/4}\int_{\pi/4}^{\pi/3} \sin 2x dx}\)

OpenStudy (anonymous):

yea that's what i did ..

OpenStudy (anonymous):

how do I evaluate without using the calculator tho?

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{1}{\pi/3-\pi/4}\int_{\pi/4}^{\pi/3} \sin 2x dx}\) \(\large \mathbb{\frac{1}{\pi/3-\pi/4} \frac{-\cos 2x}{2}\Big|_{\pi/4}^{\pi/3} }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{-6}{\pi} \cos 2x\Big|_{\pi/4}^{\pi/3} }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{-6}{\pi} [ \cos(2\pi/3) - \cos (\pi/2)] }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{-6}{\pi} [ \cos(\pi - \pi/3) - 0] }\) \(\large \mathbb{\frac{-6}{\pi} [ \cos(\pi - \pi/3) - 0] }\) \(\large \mathbb{\frac{-6}{\pi} [ -\cos( \pi/3) ] }\) \(\large \mathbb{\frac{-6}{\pi} [ -1/2 ] }\) \(\large \mathbb{\frac{3}{\pi} }\)

OpenStudy (anonymous):

the right answer is suppose to be - 1/ 6pi tho

ganeshie8 (ganeshie8):

for \(\sin(2x)\), the average value in interval \((\pi/4, \pi/3)\) is \(3/\pi\) check ur question again

OpenStudy (anonymous):

the question is correct.

ganeshie8 (ganeshie8):

if u think a bit, u wil see that average value has to be POSITIVE. wat u saying is the answer is negative.

OpenStudy (anonymous):

it is negative because the integral of sin is - cos

ganeshie8 (ganeshie8):